# Configurable photonic simulator for quantum field dynamics

### v1

@@ VI Discussion
We established the Optical Time Algorithm (OTA) for quantum-simulating the dynamics of various free scalar field theories. Our encoding is both intuitive and efficient: a single layer of phase shifters evolves the eigenmodes in time, while the physical properties of the underlying theory are encoded in two fixed beam splitter arrays followed by single-mode squeezers. In this way, a wide landscape of QFT dynamics can be quantum-simulated by adapting the parameters of a few optical elements, including (real- and complex-valued) relativistic, non-local, and non-relativistic theories as well as curved spacetimes (also exotic combinations of the latter can be envisioned), with little to no constraints on the theories’ parameters, the simulation time, or the spacetime dimension. Regarding applications, we focused on the light-cone-like spreading of quantum entanglement and correlations after a quench for varying coupling range, in which case the OTA reduces to a squeezing layer followed by passive linear optical elements. We found that $10-20$ modes are sufficient to capture field-theoretic predictions, while only a handful of those modes need to be measured.
@@
We envision four main routes for generalizations and further applications. First, in analyzing the information dynamics of quantum fields, we have focused on pure Gaussian initial states such as the vacuum, in which case the OTA attains an elementary form. This was mainly motivated by the ubiquity of Gaussian states in quantum optics experiments. However, the OTA does not rely on simple input states. Given the growing range of techniques for preparing non-Gaussian states in the lab, for instance, via conditioning schemes [136, 137], the OTA is readily applicable for studying non-Gaussian scenarios. In this context, it is of particular interest to consider the conditioning of the particle numbers at (some of) the output modes using photon-number-resolving detectors, for which the experimental setup is precisely given by the Gaussian boson sampler. This would grant access to the symmetry sectors of the (local) particle number operator, which, given a suitable readout technique, could enable the readout of symmetry-resolved entanglement entropies [20, 138].
@@
Second, while the OTA maintains its form in time-dependent scenarios whenever the diagonalization transformation (realized by the BS layers) remains time-independent, the situation becomes substantially more complex when considering general time dependencies, which are evident, for example, in the context of cosmology [27] and driven Bose-Einstein condensates [20, 42]. In this case, acting with the unitary time-evolution operator on the quadratures necessitates the inclusion of time-ordering into the exponential of the corresponding symplectic matrix transformation. This induces time dependencies in the symplectic eigenvalues, thereby complicating the disentangling of temporal contributions in this more general setting.
@@
Third, as the OTA covers all relevant free theories, a natural extension would be the incorporation of interactions. The experimental realization of strong interactions at the level of few photons is challenging since nonlinear optical effects in conventional materials are typically negligible at the associated light powers. However, in the last two decades significant advances have been made towards this goal [139], e.g., by interfacing optical modes with single atoms or atomic ensembles in cavity QED systems [140] and by using hybrid states of light and atomic media [141, 142, 143]. We speculate that the key idea of our approach, that is, isolating the time-dependence and encoding the Hamiltonian into a fixed optical circuit, is generalizable to weakly interacting theories in a perturbative fashion. Beyond the perturbative regime, a Trotterization of the dynamics might be necessary.
Fourth, our application of the OTA for studying information dynamics constitutes a proof of principle regarding the accessibility of continuum physics with a modest number of optical modes and moderate squeezing. For this and other potential applications, a more in-depth analysis of experimental constraints, such as noise and photon losses, is warranted. To this end, one may consider alternative decomposition schemes for the beam splitter layer, which can be more robust against experimental imperfections [144, 145].
@@
Recent advances towards scalable multimode photonic platforms using spatial [49, 146, 147, 148, 149], temporal [51, 150], and frequency  [151] degrees of freedom of light provide exciting opportunities for experimental realizations of the quantum field simulations proposed in this work. Although we emphasized photonic implementations due to their high level of configurability, other setups concerned with bosonic modes can also be envisioned, including circuit QED [152] and ultracold atoms [97] trapped in highly configurable optical potentials [18, 23].
@@ Acknowledgements
We thank René Sondenheimer for useful discussions. T.H. acknowledges support from the European Union under project ShoQC within the ERA-NET Cofund in Quantum Technologies (QuantERA) program, as well as from the F.R.S.- FNRS under project CHEQS within the Excellence of Science (EOS) program. M.D. and M.G. are supported by funding from the German Research Foundation (DFG) under the project identifier 398816777-SFB 1375 (NOA).
@@
We provide the code used for generating Figs. 4 and 5 together with an algorithm that computes all circuit parameters given a Hamiltonian matrix $H$ via the GitHub repository [90].
@@ A.2.2 Squeezer layer and gauge fixing
We remark that our choice of gauge fixing improves the efficiency over a direct implementation of $S_{\text{SD}}$ as in (54): choosing $G_{1}=\mathds{1}_{N}$ and $G_{2}=\mathds{O}_{N}$ (corresponding to $G=\mathds{1}_{2N}$) leads to $P_{N}$ appearing on the anti-diagonal blocks of $P$ instead. This would require an additional layer of phase shifters – which amounts to rotating $P$ back to diagonal form – before the beam splitter array.
A.2.3 Decomposing the beam splitter array
Inspired by [66], we further decompose the orthosymplectic matrix $P_{N}$ by means of a QR decomposition
@@ A.2.3 Decomposing the beam splitter array
We engineer a sequence of Givens rotations that results in a diagonal matrix where all but the last entry are one, i.e., $\mathcal{P}_{N}^{\pm}=\text{diag}(1,\dots,\pm 1)$. Denoting by $p_{jk}$ the elements of the initial matrix $P_{N}$, we choose for all $k$ such that $1\leq k<N$, as long as there is a column $j>k$ with $p_{kj}$ different from zero, the beam splitter transmittivity $\theta$ and some $r>0$ such that
@@
with Eq. (14) understood. In this way, all beam splitter parameters can be identified iteratively.
@@
The last step is to treat the two possible signs of $\mathcal{P}_{N}^{\pm}$ in Eq. (68). When the last entry is positive, we obtain $P=S^{-1}_{\text{BS}}(\theta)$, which is what we wanted to show. For the negative sign, we require an additional phase shift on the last mode $P=S_{\text{PS}}(0,\dots,\pi)S^{-1}_{\text{BS}}(\theta)$. However, this phase can be absorbed by adjusting the gauge: using that $S_{\text{PS}}(0)=S_{\text{PS}}(\pi)=-\mathds{1}_{2}$ is proportional to the identity, we can rewrite the full symplectic diagonalizing matrix as $S_{\text{SD}}=S_{\text{Sq}}(z)P=S_{\text{PS}}(0,\dots,\pi)S_{\text{Sq}}(z)S^{-1}_{\text{BS}}(\theta)$, where we used the standard gauge fixing ($G_{1}=\mathcal{O}_{N},G_{2}=\mathds{1}_{N}$). Transforming the gauge to $G\to S_{\text{PS}}^{-1}(0,\dots,\pi)G$, such that $G_{2}=\mathcal{P}_{N}^{-}$, supersedes the additional phase shift, while leaving the squeezer unchanged. It is straightforward to check that this gauge still satisfies (59), which completes our proof.
@@ Appendix C Fractional Laplacian theory on a finite lattice
In flat space, the Laplace operator $-\partial_{x}^{2}$ is diagonalized by a Fourier transform to momentum space. Hence, arbitrary powers of the Laplace operator $(-\partial_{x}^{2})^{\alpha/2}$ can be treated by raising its momentum-space eigenvalues to the same power and then transforming back to position space [106]. The eigenvalue of the term $\phi(-\partial^{2}_{x})^{\alpha/2}\phi$ is dictated by the fractional dispersion relation (36) in the massless limit $m\to 0$. Adapting this logic to the periodic lattice results in
@@
and discrete momenta labeled by $k\in\mathcal{K}=\{0,\dots,N-1\}$.
@@ D.1 Finite size effects
In a finite volume $L<\infty$ with periodic boundary conditions, quasi-particles have discrete momenta and reenter from the left (right) when traveling to the right (left) and passing through the boundary at $L$ (00). The integral over momenta in (48) is replaced by a sum and the condition $\min(2|v(p)|t,\ell)$, which can be expanded into the two terms $2|v(p)|t\times\theta(l-2|v(p)|t)+\ell\times\Theta(2|v(p)|t-l)$, has to be modified to account for reentrances of quasi-particles, see Fig. 6. Then, the Rényi-2 entanglement entropy becomes [82, 83]
@@
A similar argument can be made for adapting the Rényi-2 mutual information (49) to finite system sizes. Due to the circular geometry, the two subregions are separated by $d$ as well as the complementary distance $d^{\prime}=L-2\ell-d$ as quasi-particles contributing to the mutual information can also move in the opposite direction, yielding one additional contribution, see left vs. right panel in Fig. 6. Further, each quasi-particle pair contributes to the mutual information twice: once after having traveled a distance $(L+d)/2$, but also for $(L+d^{\prime})/2$, see the green diamonds in Fig. 6. In total, there are four contributions to mutual information before revivals, i.e., this process repeating itself. Hence, after introducing the function
Figure 6: Trajectories of quasi-particles emerging at $A_{1}$ and $A_{2}$ and moving towards each other directly (left panel) or indirectly (right panel). This results in one (three) additional finite-size contribution(s) in the Rényi-2 entanglement entropy (71) [mutual information (73)] compared to the infinite volume formulas presented in the main text. The ballistic motion is periodic, leading to momentum-dependent revivals.
@@
$\begin{split}I_{2}(t,\ell,d)=\sum_{k\in\mathcal{K}}\Big{[}&i_{2,k}(d)+i_{2,k}(d^{\prime})\\ &+i_{2,k}(L+d)+i_{2,k}(L+d^{\prime})\Big{]}.\end{split}$ (73)

### v2

@@ VII Discussion
We established the Optical Time Algorithm (OTA) for quantum-simulating the dynamics of various free scalar field theories. Our encoding is both intuitive and efficient: a single layer of phase shifters evolves the eigenmodes in time, while the physical properties of the underlying theory are encoded in two fixed interferometer arrays followed by single-mode squeezers. In this way, a wide landscape of QFT dynamics can be quantum-simulated by adapting the parameters of a few optical elements, including (real- and complex-valued) relativistic, non-local, and non-relativistic theories as well as curved spacetimes (also exotic combinations of the latter can be envisioned), with little to no constraints on the theories’ parameters, the simulation time, or the spacetime dimension. The structure of the OTA overcomes the key constraint of other decompositions: the dynamics is entirely encoded in a single phase-shift layer, while other components remain time-independent.
Regarding applications, we focused on the light-cone-like spreading of quantum entanglement and correlations after a quench for varying coupling range. We considered initial product coherent states in which case the OTA reduces to a squeezing layer followed by passive linear optical elements, resembling the setup of boson sampling experiments. We found that $10$ to $20$ modes are sufficient to capture field-theoretic predictions, also under realistic experimental constraints, while only a handful of those modes need to be measured. The tight link between the OTA and Gaussian boson sampling highlights the OTA’s prospects for addressing classically hard problems on quantum hardware.
@@
We envision four main routes for generalizations and further applications. First, in analyzing the information dynamics of quantum fields, we have focused on pure Gaussian initial states such as the vacuum, in which case the OTA attains an elementary form. This was mainly motivated by the ubiquity of Gaussian states in quantum optics experiments. However, the OTA does not rely on simple input states. Given the growing range of techniques for preparing non-Gaussian states in the lab, for instance, via conditioning schemes [154, 39, 140], the OTA is readily applicable for studying non-Gaussian scenarios. Given the inherent complexity of non-Gaussian states [101], this approach may offer advantages over classical methods even when relying on homodyne measurements. In this context, it is also of interest to consider conditioning the particle numbers at only some of the output modes using photon-number-resolving detectors, for which the experimental setup is precisely given by the Gaussian boson sampler. This would grant access to the symmetry sectors of the local particle number operator, which, given a suitable readout technique, could enable the readout of symmetry-resolved entanglement entropies in close-to-continuum settings [97, 67].
@@
Second, while the OTA maintains its form in time-dependent scenarios whenever the diagonalization transformation (realized by the interferometers) remains time-independent, the situation becomes more complex when considering general time dependencies, as is evident, for example, in driven Bose-Einstein condensates [26]. In this case, acting with the unitary time-evolution operator on the quadratures requires including time-ordering in the exponential of the corresponding symplectic matrix transformation. This induces time dependencies in the symplectic eigenvalues, thereby complicating the disentangling of temporal contributions.
@@
Third, since the OTA encompasses all relevant free theories, a natural extension is to incorporate interactions [77]. The experimental realization of strong interactions at the level of a few photons is challenging, as nonlinear optical effects in conventional materials are typically negligible at the associated light powers. However, in the last two decades, significant advances have been made towards this goal [40], e.g., by interfacing optical modes with single atoms or atomic ensembles in cavity QED systems [155] and by using hybrid states of light and atomic media [13, 119, 59]. We speculate that the key idea of our approach, that is, isolating the time-dependence and encoding the Hamiltonian into a fixed optical circuit, is generalizable to weakly interacting theories in a perturbative fashion. Beyond the perturbative regime, a Trotterization of the dynamics might be necessary, in which case the OTA would form the elementary building block of the non-interacting part.
@@
Fourth, recent advances towards scalable multimode photonic platforms using spatial [157, 165, 28, 66, 158], temporal [98, 27], and frequency [60] degrees of freedom of light provide exciting opportunities for experimental realizations of the quantum field simulations proposed in this work. Although we emphasized photonic implementations due to their high configurability, other setups involving bosonic modes can also be envisioned, including circuit QED [21] and ultracold atoms [121] trapped in highly configurable optical potentials [70, 133].
@@ Acknowledgements
We thank Benjamin Brecht, Dario Cilluffo, Martin Plenio, Christine Silberhorn, René Sondenheimer, and Michael Stefzsky for useful discussions. We thank Alexander Naumann and Robin Strahlendorf for providing a Python package for simulating photon loss [107]. T.H. acknowledges support from the European Union under project ShoQC within the ERA-NET Cofund in Quantum Technologies (QuantERA) program, the F.R.S.-FNRS under project CHEQS within the Excellence of Science (EOS) program, the BMBF project PhoQuant (Grant No. 13N16110), and the EU project SPINUS (Grant No. 101135699). M.D. and M.G. are supported by funding from the German Research Foundation (DFG) under the project identifier 398816777-SFB 1375 (NOA).
@@
We provide the code used for generating Figs. 47 and 10 together with an algorithm that computes all circuit parameters given a Hamiltonian matrix $H$ via the GitHub repository [47]. The code for loss simulation is publicly accessible via GitHub [107].
@@ A.2.2 Squeezer layer and gauge fixing
We remark that our choice of gauge fixing improves the efficiency over a direct implementation of $S_{\text{SD}}$ as in (59): choosing $G_{1}=\mathds{1}_{N}$ and $G_{2}=\mathds{O}_{N}$ (corresponding to $G=\mathds{1}_{2N}$) leads to $P_{N}$ appearing on the anti-diagonal blocks of $P$ instead. This would require an additional layer of phase shifters—which amounts to rotating $P$ back to diagonal form—before the interferometer array.
A.2.3 Decomposing the interferometer
A variety of schemes can be implemented to decompose the interferometers, e.g. Reck’s [125] and Clements’ [45], see [44] for an overview. Inspired by the first method, as visible in the zoom-in in Fig. 8, we further decompose the orthosymplectic matrix $P_{N}$ by means of a QR decomposition
@@ A.2.3 Decomposing the interferometer
Figure 8: Pictorial representation of the OTA for $N=5$. The interferometer layers are decomposed here according to a Reck-inspired method.
We engineer a sequence of Givens rotations that results in a diagonal matrix where all but the last entry are one, i.e., $\mathcal{P}_{N}^{\pm}=\text{diag}(1,\dots,\pm 1)$. Denoting by $p_{jk}$ the elements of the initial matrix $P_{N}$, we choose for all $k$ such that $1\leq k<N$, as long as there is a column $j>k$ with $p_{kj}$ different from zero, the beam splitter transmissivity $\theta$ and some $r>0$ such that
@@
with the triangular ordering, compatible with Fig. 8,
$\begin{split}S_{\text{I}}&\equiv S_{\text{BS}}(\theta_{12})S_{\text{BS}}(\theta_{13})\dots S_{\text{BS}}(\theta_{1N})\\ &\hskip 39.26494pt\times S_{\text{BS}}(\theta_{23})\dots S_{\text{BS}}(\theta_{2N})\\ &\hskip 61.17325pt\times\dots\\ &\hskip 79.09862pt\times S_{\text{BS}}(\theta_{N-1N}),\end{split}$ (74)
understood. In this way, all beam splitter parameters can be identified iteratively.
@@
The last step is to treat the two possible signs of $\mathcal{P}_{N}^{\pm}$ in Eq. (73). When the last entry is positive, we obtain $P=S^{-1}_{\text{I}}$, which is what we wanted to show. For the negative sign, we require an additional phase shift on the last mode $P=S_{\text{PS}}(0,\dots,\pi)S^{-1}_{\text{I}}$. However, this phase can be absorbed by adjusting the gauge: using that $S_{\text{PS}}(0)=S_{\text{PS}}(\pi)=-\mathds{1}_{2}$ is proportional to the identity, we can rewrite the full symplectic diagonalizing matrix as $S_{\text{SD}}=S_{\text{Sq}}(z)P=S_{\text{PS}}(0,\dots,\pi)S_{\text{Sq}}(z)S^{-1}_{\text{I}}$, where we used the standard gauge fixing ($G_{1}=\mathcal{O}_{N},G_{2}=\mathds{1}_{N}$). Transforming the gauge to $G\to S_{\text{PS}}^{-1}(0,\dots,\pi)G$, such that $G_{2}=\mathcal{P}_{N}^{-}$, supersedes the additional phase shift, while leaving the squeezer unchanged. It is straightforward to check that this gauge still satisfies (64), which completes our proof.
A.3 Time-dependent case
We generalize the OTA to time-dependent Hamiltonian matrices $H=H(t)$ for time-independent eigenvectors $P\neq P(t)$, which ensures $\bm{H}(t_{1}),\bm{H}(t_{2})]=0$. Starting once again from Williamson’s theorem (8), i.e., $H(t)=S_{\text{SD}}^{\intercal}(t)D(t)S_{\text{SD}}(t)$, and using the polar decomposition $S_{\text{SD}}=S_{\text{Sq}}(t)P$ [see Eqs. (65) and (67)] results in
$\bm{H}(t)=\dfrac{1}{2}\bm{U}_{\text{I}}\bm{r}^{\intercal}S_{\text{Sq}}^{\intercal}(t)D(t)S_{\text{Sq}}(t)\bm{r}\bm{U}_{\text{I}}^{\dagger},$ (75)
with $P=S_{\text{I}}^{-1}$ and $S_{\text{I}}\bm{r}=\bm{U}_{\text{I}}^{\dagger}\bm{r}\bm{U}_{\text{I}}$ understood. The corresponding unitary time evolution operator becomes
$\begin{split}\bm{U}(t)&=\exp\left[-i\int_{0}^{t}\mathrm{d}t^{\prime}\bm{H}(t^{\prime})\right]\\ &=\bm{U}_{\text{I}}\exp\left[-\frac{i}{2}\bm{r}^{\intercal}\int_{0}^{t}\mathrm{d}t^{\prime}S_{\text{Sq}}^{\intercal}(t^{\prime})D(t^{\prime})S_{\text{Sq}}(t^{\prime})\bm{r}\right]\bm{U}_{\text{I}}^{\dagger}.\end{split}$ (76)
As $\bm{U}_{\text{I}}$ diagonalizes the unitary, the remaining integral runs over diagonal matrices only. Hence, we recover the OTA when introducing time-integrated eigenvalues
$\tilde{\lambda}_{j}^{\phi(\pi)}=\int_{0}^{t}\mathrm{d}t^{\prime}\lambda_{j}^{\phi(\pi)}(t^{\prime}),$ (77)
which leads to
$S(t)=S_{\text{I}}\,S_{\text{Sq}}^{-1}[z(t)]\,S_{\text{PS}}[\varphi(t)]\,S_{\text{Sq}}[z(t)]\,S_{\text{I}}^{-1}.$ (78)
The single-mode optical parameters become time-dependent, i.e., $\varphi_{j}(t)=\tilde{d}_{j}(t),z_{j}(t)=-(1/2)\ln\tilde{\gamma}_{j}(t)$, with Eqs. (17) and Eqs. (18) generalizing to $\tilde{\gamma}_{j}^{2}(t)=\tilde{\lambda}^{\phi}_{j}(t)/\tilde{\lambda}^{\pi}_{j}(t)$ and $\tilde{d}_{j}^{2}(t)=\tilde{\lambda}^{\phi}_{j}(t)\tilde{\lambda}^{\pi}_{j}(t)$, respectively.
A.4 Coherent inputs
[…]
The main simplification for coherent inputs hinges on the fact that the single-mode (SM) part of the OTA, i.e.,
$S_{\text{SM}}(z,\varphi)=S_{\text{Sq}}^{-1}(z)S_{\text{PS}}(\varphi)S_{\text{Sq}}(z),$ (79)
can be rewritten as a complex squeezer, see Eq. (24). To show this, we must enforce consistency of the quantum state’s first and second moments only at the single-mode level. The latter implies the condition
$(S_{\text{SM}}S_{\text{SM}}^{\intercal})(z,\varphi)=(S_{\text{CSq}}S_{\text{CSq}}^{\intercal})(\zeta),$ (80)
with $\zeta=\xi e^{i\Phi}$ understood, which is equivalent to the non-linear system of equations
$\begin{split}\cosh(2\xi)&=\cos^{2}(\varphi)+\sin^{2}(\varphi)\cosh(4z),\\ \sin^{2}(\varphi)\sinh(4z)&=-\cos(\Phi)\sinh(2\xi),\\ \sin(2\varphi)\sinh(2z)&=-\sin(\Phi)\sinh(2\xi).\end{split}$ (81)
While the first equation amounts to the first equation in (25), dividing the third by the second gives
$\tan\Phi=\dfrac{\cot\varphi}{\cosh(2z)},$ (82)
thus proving the second equation in (25) as well. This also completes the proof of (26) for vacuum inputs, as the first-moment condition turns trivial.
For coherent inputs, it remains to identify the mapping of the coherent amplitudes. From a geometrical perspective in phase space, we need to displace the modified initial state with amplitude $\alpha^{\prime}$ yet another time when matching $S_{\text{SM}}$ with a complex squeezer. To show this, we switch to unitary operators and employ several commutation relations. First, we note that
[…]
where we used the elementary commutation relations
$\begin{split}\bm{U}_{\text{Sq}}(z)\bm{D}(\alpha^{\prime})&=\bm{D}(\alpha^{\prime}\cosh z-\alpha^{\prime*}\sinh z)\bm{U}_{\text{Sq}}(z),\\ \bm{U}_{\text{PS}}(\varphi)\bm{D}(\alpha^{\prime})&=\bm{D}(\alpha^{\prime}e^{-i\varphi})\bm{U}_{\text{PS}}(\varphi),\end{split}$ (84)
leading to $\alpha^{\prime\prime}$ depending on $\alpha^{\prime},z$ and $\varphi$ through
$\alpha^{\prime\prime}=\alpha^{\prime}\cos\varphi-i\sin\varphi[\alpha^{\prime}\cosh(2z)-\alpha^{\prime*}\sinh(2z)].$ (85)
[…]
$\alpha^{\prime\prime\prime}=\alpha^{\prime\prime}\cosh\xi+\alpha^{\prime\prime*}e^{i\Phi}\sinh\xi,$ (86)
thereby completing the proof.
@@ Appendix C Fractional Laplacian theory on a finite lattice
In flat space, the Laplace operator $\partial_{x}^{2}$ is diagonalized by a Fourier transform to momentum space. Hence, arbitrary powers of the Laplace operator $(-\partial_{x}^{2})^{\alpha/2}$ can be treated by raising its momentum-space eigenvalues to the same power and then transforming back to position space [95]. The eigenvalue of the term $\phi(-\partial^{2}_{x})^{\alpha/2}\phi$ is dictated by the fractional dispersion relation (37) in the massless limit $m\to 0$. Adapting this logic to the periodic lattice results in
@@
and discrete momenta labeled by $k\in\mathcal{K}=\{0,\dots,N-1\}$. Note here that discretization can become ambiguous for other than periodic boundary conditions [95].
@@ D.1 Finite size effects
In a finite volume $L<\infty$, quasi-particles have discrete momenta and their dynamics are influenced by boundary conditions. We discuss the two most prominent choices: periodic and open boundaries.
Figure 9: Boundary effects relevant for identifying the finite-size Rényi-2 mutual information. We trace all trajectories of two quasi-particles emerging at $A_{1}$ and $A_{2}$ that contribute to the buildup of correlations—marked by crossing trajectories. (a) Periodic boundary conditions. Parallel trajectories never cross. Hence, only particle pairs moving towards each other directly (left panel) or indirectly (right panel) contribute. This results in one (three) additional finite-size contribution(s) to the Rényi-2 entanglement entropy (89) [mutual information (91)] compared to the infinite volume formulas presented in the main text. As the ballistic motion is periodic, we will observe momentum-dependent revivals. (b) Open boundary conditions. In addition to the previous two possibilities, quasi-particles moving in parallel now do contribute to correlations (two rightmost panels). Further, the periodicity doubles.
D.1.1 Periodic boundary conditions (PBCs)
For periodic boundaries, quasi-particles reenter from the left (right) when traveling to the right (left) and passing through the boundary at $L$ (0). The integral over momenta in (52) is replaced by a sum and the condition $\min(2|v(p)|t,\ell)$, which can be expanded into the two terms $2|v(p)|t\times\theta(l-2|v(p)|t)+\ell\times\Theta(2|v(p)|t-l)$, has to be modified to account for reentrances, see Fig. 9(a). Then, the Rényi-2 entanglement entropy becomes [38, 103]
@@ D.1.1 Periodic boundary conditions (PBCs)
A similar argument can be made for adapting the Rényi-2 mutual information (53) to finite system sizes. Due to the circular geometry, the two subregions are separated by $d$ as well as the complementary distance $d^{\prime}=L-2\ell-d$ as quasi-particles contributing to the mutual information can also move in opposite directions, yielding one additional contribution, see left vs. right panel in Fig. 9(a). Further, each quasi-particle pair contributes twice to the mutual information: once after having traveled a distance $(L+d)/2$, but also for $(L+d^{\prime})/2$, see the green diamonds in Fig. 9(a). In total, there are four contributions to mutual information before revivals, i.e., this process repeating itself. Hence, after introducing the function
@@
$\begin{split}I_{2}(t,\ell,d)=\sum_{k\in\mathcal{K}}\Big[&i_{2,k}(d)+i_{2,k}(d^{\prime})\\[-8.61108pt] &+i_{2,k}(L+d)+i_{2,k}(L+d^{\prime})\Big].\end{split}$ (91)
Note that $\ell$ appears in the previous equation: Fig. 9 sketches $M=1$, while we consider $|A_{1}|=|A_{2}|=M>1$.
D.1.2 Open boundary conditions (OBCs)
For open boundaries, quasi-particles reaching the boundary are reflected, corresponding to a velocity inversion $v_{k}\rightarrow-v_{k}$. Whether the particles’ phase jumps during elastic scattering or not is specified by Dirichlet (fixed ends $\phi_{0}=\phi_{N}=0$, phase jump of $\pi$) or von Neumann (free ends $(\phi_{1}-\phi_{0})/\epsilon=(\phi_{N}-\phi_{N-1})/\epsilon=0$, no phase jump) boundary conditions. For our analysis, however, phase jumps are irrelevant.
When considering the relativistic theory (30) and von Neumann boundaries, the Hamiltonian matrix given in Eq. (32) modifies to
$H_{\text{R}}^{\phi}=\begin{pmatrix}\mu+\nu&\nu&\ldots&0&0\\ \nu&\mu&\ldots&0&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&\ldots&\mu&\nu\\ 0&0&\ldots&\nu&\mu+\nu\end{pmatrix}$, (92)
with $\mu=\epsilon m^{2}-2\nu$ and $\nu=-1/\epsilon$ as before. The Hamiltonian is diagonalized by an orthogonal transformation resulting in the dispersion
$\omega^{2}_{k}=m^{2}+\dfrac{4}{\epsilon^{2}}\sin^{2}\left[\tfrac{\pi(k-1)}{2N}\right],$ (93)
which differs from the PBC result (33) by a factor of $1/2$ in the argument of the sine.
Also for open boundaries, the quasi-particle picture offers a reasonable description for how information spreads throughout the system. The discrete group velocity $v_{k}=\sin[\pi(k-1)/N]/(\epsilon\omega_{k})$, although being of slightly different functional form, is almost identical to the PBC case for $N\gtrsim 10$. Hence, $v_{\text{max}}$ is unaltered. What changes are the quasi-particle trajectories, which impact the function $g_{k}$ entering (89) and controlling the transition in the Rényi-2 entropy from linear growth (early times) to saturation (late times). For OBC, quasi-particles produced at distances $\ell/2$ from the entangling region cannot enter the region $B$ by traversing the boundary, resulting in the replacement $g_{k}=\{2|v_{k}|t/L\}\rightarrow\{|v_{k}|t/L\}$, while (89) remains intact. Therefore, the time until saturation doubles relative to PBC, i.e., $2\tau/N$, which translates into a doubled revival time $L/v_{\text{max}}$, as shown in Fig. 10(a).
For the Rényi-2 mutual information, we must account for multiple factors. Importantly, in this case, the function $g_{k}$ is the same one used in rthe PBC case. This follows after comparing the green diamonds in the left diagrams of Fig. 9(a) and Fig. 9(b)—the time required for the first build-up of correlations (and thus also the revival time) is identical. However, the number of possible trajectories entering the finite-size mutual information formula increases from four, as in Eq. (91), to eight. Quasi-particles initially moving in parallel contribute to overall correlations after being reflected at the boundaries, see, e.g., the rightmost panel in Fig. 9(b). Moreover, the complementary distance $d^{\prime}=L-2\ell-d$ considered for PBC, see left panel in Fig. 9(a), is not unique anymore. It must be replaced by $d_{\text{L}}$ and $d_{\text{R}}$, which denote the distances from the left region to the left wall and from the right region to the right wall, respectively, as shown in the leftmost panel of Fig. 9(b). For given $d_{\text{L}}$, one can compute $d_{\text{R}}$ via $d_{\text{R}}=L-2\ell-d-d_{\text{L}}$. In summary, the finite-size mutual information for open boundaries attains the form
$\begin{split}&I_{2}(t,\ell,d,d_{\text{L}})\\ &=\sum_{k\in\mathcal{K}}\Big[i_{2,k}(d)+i_{2,k}(L+d_{\text{L}}+d_{\text{R}})\\[-9.47217pt] &\hskip 34.14322pt+i_{2,k}(2L+d)+i_{2,k}(3L+d_{\text{L}}+d_{\text{R}})\\ &\hskip 34.14322pt+i_{2,k}(2d_{\text{L}}+d+\ell)+i_{2,k}(2d_{\text{R}}+d+\ell)\\ &\hskip 34.14322pt+i_{2,k}(2L+2d_{\text{L}}+d+\ell)\\[-4.73611pt] &\hskip 34.14322pt+i_{2,k}(2L+2d_{\text{R}}+d+\ell)\Big],\end{split}$ (94)
with $i_{2,k}(x)$ as in (90).
In Fig. 10, we show the quench dynamics of the relativistic scalar as discussed in Sec. V.2 with the same parameters as in Fig. 4, but open (von Neumann) instead of periodic boundaries. The early-time dynamics of the mutual information between two separated spatial regions essentially agree for periodic and open boundaries, i.e., $I_{2}^{\text{OBC}}(t,\ell,d)\simeq I_{2}^{\text{PBC}}(t,\ell,d)$ [compare Fig. 10(b) with Fig. 4(b)], which we trace back to their similar first quasi-particle crossing times. Note, however, that their late-time evolution can differ. In contrast, the Rényi-2 entropy for OBC evolves at roughly half the pace of PBC, i.e., $S_{2}^{\text{OBC}}(t,\ell)\simeq S_{2}^{\text{PBC}}(t/2,\ell)$ [Fig. 10(a) vs. Fig. 4(a)], since quasi-particles moving in opposite directions can no longer shortcut via periodic boundaries, which delays their crossings by a factor of two. As for our observations in Sec. V.2, the photonic simulations are in notable agreement with QFT predictions for similar system sizes $N\gtrsim 10-20$.
Figure 10: Same analysis as in Fig. 4 for open boundaries (von Neumann) breaking circulanticity. (a) The Rényi-2 entanglement entropy builds up more slowly, with a linear increase until $2\tau/N=\ell/(Nv_{\text{max}})\approx 0.98$ set by the fastest quasi-particles. (b) The Rényi-2 mutual information remains largely unaltered, especially for larger mode counts $N\gtrsim 20$.